Please help!
I simplified (3+i)^2 and got 9-1 but it was wrong.
thanks

Answer by Alan3354(69443)

(3+i)^2
= (3+i)*(3+i)
= 9 + 6i + i^2
= 9 + 6i - 1
= 8 + 6i